OpenStudy (dumbsearch2):

The yogurt shop wants to mix two toppings together to create a new topping that will sell for $7 per pound. The caramel crunch costs $3 per pound, and pecan turtles cost $8 dollars per pound. How many pounds of each item must be mixed together to get 10 pounds of this new topping? (Note I'm not kidding I have a real Amazon coupon code but this may sound ridiculous but if you respond with the answer I will give you one of my 4 dollar coupons. For real. :) :))

OpenStudy (anonymous):

3c + 8p = 7 * 10 c + p = 10 3c + 8p = 70 -3c - 3p = -30 5p = 40 p = 8 lbs c = 2 lbs

OpenStudy (anonymous):

Let p be the pecan in pounds and c be the caramel crunch in pounds.

OpenStudy (anonymous):

You would make the system of equations and multiply the second equation by 3 to eliminate the x's

OpenStudy (anonymous):

@dumbsearch2 got it?

OpenStudy (dumbsearch2):

Thanks SO MUHC!!! Thanks here are the coupons for amazon: D5Q6-FBPXBK-FV6U and JAU3-ZDGGT7-MSXJ and NJYX-FU2XU8-ZAFB !!!! :)